Combining the Rules to Solve Equations
Solve for x and check your result
6x-1/5 - 2x/3=3 (fractions)

Answer by praseenakos@yahoo.com(507)





6x-2x/3= 3+1/5
\ (18x-2x)/3= (15+1)/5
16x/3 = 16/5
x= (16).(3)/(16).(5)
x = 3/5

check
substituting the value of x
6(3/5)-1/5 - 2.3/3.5
18/5- 1/5 -2/5
(18-1-2)/5
15/5= 3
therefore L.H.S = R.H.S